# PARCO: Learning Parallel Autoregressive Policies for Efficient Multi-Agent Combinatorial Optimization

### v1

@@ B.4.1 Hardware
Experiments were carried out on a machine equipped with two Intel(R) Xeon(R) Gold 6338 CPU @ 2.00GHz CPUs with a total $128$ threads and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM. During inference, we employ only one CPU and a single GPU.
@@ B.4.2 Software
We used Python 3.12, PyTorch 2.4 (Paszke et al., 2019) coupled with PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) with most code based on the RL4CO library (Berto et al., 2024a). The main OS is Ubuntu 22.04.4 LTS. There were no major difficulties when trying the code on other platforms, so we expect experiments to be reproducible in other architectures.
@@ B.5 Source Code
We value open reproducibility. We provide the source code to reproduce our experiments on Github at https://github.com/ai4co/parco.
Appendix C Additional Discussion
Following Section 6, we outline further discussion here. We additionally remark that several areas of multi-agent planning, including multi-agent pathfinding (Lin and Ma, 2023; Wang et al., 2023; Tang et al., 2024a) and social robot navigation (Li et al., 2024a; He et al., 2024) employ communication as a tool with great importance, from which we also borrow ideas such as prioritized conflict resolution (Tang et al., 2024b).
An exciting future direction going forward is the development of foundation models, in which PARCO could aid in improving the efficiency and performance in solving multi-agent CO problems. Recent examples in the literature have demonstrated the possibility of scaling up solvers with (self)supervision (Drakulic et al., 2023; Luo et al., 2023; Corsini et al., 2024; Pirnay and Grimm, 2024; Luo et al., 2024), developing multi-task solvers for VRPs (Liu et al., 2024a; Lin et al., 2024; Zhou et al., 2024; Berto et al., 2024b) as well as more general CO tasks (Drakulic et al., 2024; Jiang et al., 2024). Such approaches model CO problems as sequential autoregressive decision-making; parallelizing this process with PARCO could enrich solution speed/quality. We note that most NCO approaches in e.g., VRPs, mostly focus on solving without explicitly modeling (heterogeneous) agents as in PARCO, which is of practical relevance in several real-world tasks. Better heuristics for prioritized planning might also be found utilizing automated discovery via recent LLM approaches for CO (Liu et al., 2024b; Ye et al., 2024b). Finally, integrating PARCO in online optimization (Hottung et al., 2021; Choo et al., 2022) and end-to-end pipelines for construction and improvement, as done recently in (Berto et al., 2024a; Kong et al., 2024), would further improve the performance of PARCO given larger computational time budgets.

### v2

@@ B.4.1 Hardware
We experiment on a workstation equipped with 2 Intel(R) Xeon(R) Gold 6338 CPUs and $8$ NVIDIA RTX 4090 graphic cards with $24$ GB of VRAM each. Training runs of PARCO take less than 24 hours each. During inference, we employ only one CPU and a single GPU.
@@ B.4.2 Software
We used Python 3.12, PyTorch 2.5 (Paszke et al., 2019) coupled with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) with most code based on the RL4CO library (Berto et al., 2024a). The operating system is Ubuntu 24.04 LTS.
@@ B.5 Source Code
B.5.1 Licenses
We provide the summary of licenses Table 3.
Table 3: Summary of licenses for used assets.
Resource Type License
OR-Tools (Furnon & Perron, 2024) Code Apache License, Version 2.0
AM (Kool et al., 2019a) Code MIT License
ET (Son et al., 2024) Code Available on Github
DPN (Zheng et al., 2024a) Code MIT License
DRLLi (Li et al., 2022) Code Available on Github
2D-Ptr (Liu et al., 2024b) Code/Dataset Available on Github
HAM (Li et al., 2021a) Code MIT License
Gurobi (Gurobi Optimization, LLC, 2024) Code Commercial license (free for academic use)
MatNet (Kwon et al., 2021) Code/Dataset MIT License
RL4CO (Berto et al., 2024a) Library MIT License
B.5.2 Code Release
We value open reproducibility and release our source code at https://github.com/ai4co/parco.